# Forward modeling fluctuations in the DESI LRGs target sample using image simulations

### v1

@@ 7.4 Other secondary effects
We have identified some other effects that would potentially affect our ability to properly simulate the expected target density fluctuations in the real data. We do not see an evident mismatch when not accounting for these effects. However, we list these effects and potential treatments below.
Shape noise: Real galaxies are not perfect model galaxies. Since our injected galaxies are accurately described by the simple galaxy models, they would be slightly different from what we see in real galaxies. For the precision that we are worried about now, this effect is negligible. However, this would be a bigger issue when trying to study galaxies’ shape response, e.g., as done for weak lensing measurements [80], and the studies of magnification [81]. Realistic galaxy stamps will be needed in this analysis.
Sagittarius stream [82]: No noticeable trend due to the Sagittarius stream is present in the LRGs. The contamination from stars is extremely low for LRGs, so stars do not provide direct contamination. However, the Sagittarius stream, along with other stellar streams or Milky Way stars, can enter contamination by altering the background level of the image. As stellar streams originate from satellite galaxies, they contain many faint stars that are below the detection limit and are merged into the background. When falsely subtracting background noise that is faint Sagittarius stream stars, we may get a shifted flux measurement.
Variation in extinction coefficients [83]: The extinction coefficient denotes the extent of the impact that extinction has on each band. It is computed with stellar spectra. However, this coefficient is related to the wavelength of observation. This means that if the galaxy we observed is very different from stellar spectra, the coefficient we assume would not be accurate.
@@ 7.5 Discussion on imaging systematics mitigation
Currently, studies on imaging systematics mitigation typically employ an approach based on survey property maps. The observed footprint for a survey is separated into small pixels. Each pixel is associated with a collection of quantities like depth, PSF, $E(B-V)$, airmass, etc. Thus, we have pixelized maps for each of these quantities. Studies on imaging systematics mitigation typically find a relationship between these survey property maps and the false fluctuations of galaxy densities. Then these studies derive “imaging systematics weights”. The weights minimize the systematics trend on all these survey property maps. One weight is assigned to one pixel so that the weights form an imaging-systematics-weight map with the same pixel resolution. These kinds of methods rely on two assumptions that are not entirely true:
  • Assumption 1: No errors are present in these survey property maps.
@@
  • Assumption 2: For sources at the same location, the imaging systematics associated with them are the same, and can be represented as a number that we call “imaging systematics weight”.
@@
Our study with galactic extinction shows flaws on Assumption 1. The SFD $E(B-V)$ map is commonly used as one of the survey property maps. Therefore, such methods have the assumption that the SFD $E(B-V)$ map is perfectly accurate. After applying systematics weight, the sample tends to barely have any trend against the SFD $E(B-V)$ map. Our study challenges this approach of treating the SFD $E(B-V)$ map. Our findings imply that a portion of the observed trend within the SFD $E(B-V)$ map may be attributable to contamination stemming from CIB or other Large-Scale-Structure systematics. In section 7.2, factor 1 could result in false fluctuations of galaxies that need to be corrected with weights. Since target selection defines color cut to select LRGs, this shift is most evident around the color selection boundaries. This effect needs to be carefully corrected since its impact on different magnitudes of LRGs may vary significantly. Factor 2 is purely due to false binning and should not be corrected at all. Ideally, systematics-weighted galaxies should still have a positive trend against the SFD $E(B-V)$ map as we should not correct the trend that stems from factor 2.
@@
Our study on galaxy brightness in section 7.1 and psfsize in section 7.3 challenges Assumption 2. We show that the imaging systematics trend for DESI LRGs changes significantly when we apply a cut on $z$-band magnitude. The mechanism behind this is well-understood: Faint galaxies have more photometric scattering than bright galaxies so they are affected by the variation of galaxy depth and seeing (psfsize) to a larger extent. If two galaxies of different magnitudes have the same configuration in a survey property map, they could be affected by imaging systematics differently. However, in the current survey-property-map-based approach, they will be assigned the same “imaging systematics weight”. Applying the same correction for galaxies with different brightness will result in an over-correction in brighter galaxies and an under-correction in fainter galaxies. In both cases, they introduce an artificial signal when computing the sample’s power spectrum.
@@
One way to mitigate this issue is to separate galaxy samples into subsamples with different magnitude ranges. Then we compute imaging systematics weights independently for each subsample. However, with a lower density for each subsample, the computation of imaging systematics weight would be prone to fit the noise of each subsample. Thus, taking such a step needs to be carefully validated.
@@
Another way to deal with the issue is through forward modeling tools like Obiwan. Forward modeling can produce randoms with detailed scattering features for each source at any location. We can use such randoms to correct for magnitude-dependent imaging systematics. The randoms produced from forward modelling can predict density fluctuations at any specific color and brightness. Compared with making such predictions with real data, the advantage of predicting this density fluctuation with forward modelling randoms like Obiwan LRGs is that these randoms can reach a much larger number density and does not have true clustering signal in it. These advantages would help making the predictions precise and free from overfitting the true clustering signal.
Due to the scope of this work, we did not do actual measurements. Current Obiwan LRGs does not have a large enough density to be used as magnitude-dependent randoms. The current procedure of Obiwan is computationally expensive. However, this is still feasible with image simulations that are significantly faster than Obiwan. One idea that allows significant speed-up is to measure multiple synthetic galaxies at one location repeatedly. This would avoid repeating time-consuming stages like reading images, detecting sources, etc, and focus on the stage of source fitting. Furthermore, the procedure of source fitting can be accelerated by operating the code on GPUs, as it involves intensive utilization of matrix operations.
Another practical way to address this problem is to develop a Monte Carlo procedure to analytically estimate how galaxies are scattered under certain conditions. This procedure numerically estimates how galaxy magnitudes are altered by estimating errors generated by each survey property. It is much faster than image simulation. Such a method would need to be calibrated on image simulation to validate the accuracy of its numerical estimations and the Obiwan results we have presented could be used for such calibration.
@@ 8 Conclusions
  • We find shape bias in the COSMOS Deep catalog. We develop a procedure to eliminate this shape bias successfully, see Section 5.1.
@@
  • We see imprints of Large-Scale-Structure systematics (e.g. CIB contamination) in extinction maps based on dust emission. We find significant trends between the DESI LRGs density on the sky and the SFD $E(B-V)$ map [65], which are not predicted by Obiwan. These trends change significantly when we test against alternative $E(B-V)$ maps and are nearly removed when compared to the stellar-reddening-based $E(B-V)$ map [69]. Since CIB is correlated with galaxy clustering, we caution against any analyses of DESI LRGs eliminating this trend, as this would result in removing real large-scale structures. Meanwhile, some extent of the trends with $E(B-V)$ is real and related to the impact on the imaging depth (photometric scattering), and is predicted by our forward modeling.
  • We find that Obiwan predicts the trends with psfdepth that we observe in the real data, which are similarly strong in each of the optical imaging bands. While the overall trends are predicted, statistically, they are not consistent in the SGC. The slight mismatch we see is due to imperfections in the truth input, which has a larger impact on the deepest data in the SGC. The mismatch in the color and magnitude distribution of Obiwan LRGs and DESI LRGs is the main source of the systematics trend deviation in this region. We further demonstrate that this issue can be greatly mitigated by applying a ‘color matching’ weight to Obiwan LRGs to make its color distribution match with that of DESI LRGs. The details are discussed in section 7.1.
  • We find that the trends with imaging depth depend strongly on the $z$-band magnitude. When selecting the data to have $z<20.5$, only  1% level variations are found in the observed data with the imaging depth and Obiwan predicts no significant trend. However, when selection $z>20.5$, up to 20% trends are observed in the real data and predicted by Obiwan. This suggests it is important to properly model the color/brightness dependence of systematic trends in the DESI LRGs target sample when producing any corrections for use in galaxy clustering measurements. Forward modeling randoms like DESI LRGs can be a powerful catalog to correct for imaging systematics that varies with intrinsic galaxy brightness.
  • We find signs of non-simulated density variation against psfsize, which is especially prominent for extended galaxies. This is likely due to not taking into account PSF error and/or background subtraction error by Obiwan. This effect results in LRGs around color selection boundaries having different properties. The effect is not very noticeable in the overall number density of LRGs, but it needs to be properly taken into account in cosmological analysis.
@@
  • By comapring the systematics trends in real galaxies, and the synthetic galaxies from forward modeling tools like Obiwan, we can identify signs of unknown imaging systematics, and investigate relevant resources to figure out the root cause of them. Replicating imaging systematics in real galaxies through forward modeling tools like Obiwan is challenging work. A perfect match between real and synthetic galaxies is only possible when all potential contaminations are properly considered in the image simulation settings. However, a mismatch of density fluctuations in the synthetic galaxies also provides us valuable insights into what unknown systematics could possibly be missed. Identifying missing factors is vital for proper imaging systematics mitigation.
@@
We produce the Obiwan LRGs catalog that traces imprints of DESI LRGs in the real images and can be used for future studies. The product is currently available internally for DESI members  [^10] https://data.desi.lbl.gov/desi/survey/catalogs/image_simulations/LRG/. The data will be uploaded to zenodo before publication [^11] https://zenodo.org/uploads/11260267. In the future, the catalogs used in this analysis will be made public along with the Data Release 1 121212 https://data.desi.lbl.gov/doc/releases/.
@@ Acknowledgements
We thank Alex Krolewski for reading and providing feedback for this work.

### v2

@@ 7.4 Other secondary effects
We have identified other effects that would potentially affect our ability to properly simulate the expected target density fluctuations in the real data. We do not see an evident mismatch when these effects are not being accounted for. However, we list these effects and potential treatments below.
Shape noise: Real galaxies are not perfect model galaxies. Since our injected galaxies are accurately described by the simple galaxy models, they would be slightly different from what we see in real galaxies. This effect is negligible for the precision we are worried about now. However, this would be a more significant issue when trying to study galaxies’ shape response, e.g., as done for weak lensing measurements [87], and the studies of magnification [88]. Realistic galaxy stamps will be needed in this analysis.
Sagittarius stream [89]: No noticeable trend due to the Sagittarius stream being present in the LRGs. The contamination level from stars is extremely low for LRGs, so stars do not provide direct contamination. However, the Sagittarius stream, along with other stellar streams or Milky Way stars, can enter contamination by altering the background level of the image. As stellar streams originate from satellite galaxies, they contain many faint stars below the detection limit and merge into the background. We may get a shifted flux measurement when falsely subtracting background noise that is faint Sagittarius stream stars.
Variation in extinction coefficients [73]: The extinction coefficient denotes the extent of extinction’s impact on each band. It is computed with stellar spectra. However, this coefficient is related to the wavelength of observation. This means that if the galaxy we observed differs from stellar spectra, the coefficient we assume would not be accurate.
@@ 7.5 Discussion on imaging systematics mitigation
Currently, studies on imaging systematics mitigation typically employ an approach based on survey property maps. The observed footprint for a survey is separated into small pixels. Each pixel is associated with a collection of quantities like depth, PSF, $E(B-V)$, airmass, etc. Thus, we have pixelized maps for each of these quantities. Studies on imaging systematics mitigation typically find a relationship between these survey property maps and the false fluctuations of galaxy densities. Then, these studies derive “imaging systematics weights”. The weights minimize the systematics trend on all these survey property maps. One weight is assigned to one pixel so that the weights form an imaging-systematics-weight map with the same pixel resolution. These kinds of methods rely on two assumptions that are not entirely true:
  • Assumption 1: No errors exist in these survey property maps.
@@
  • Assumption 2: For sources at the same location, the imaging systematics associated with them are the same and can be represented as a number that we call “imaging systematics weight”.
@@
Our study with galactic extinction shows flaws on Assumption 1. The SFD $E(B-V)$ map is commonly used as a survey property map. Therefore, such methods have the assumption that the SFD $E(B-V)$ map is perfectly accurate. After applying systematics weight, the sample tends to barely have any trend against the SFD $E(B-V)$ map. Our study challenges this approach of treating the SFD $E(B-V)$ map. Our findings imply that a portion of the observed trend within the SFD $E(B-V)$ map may be attributable to contamination stemming from CIB or other Large-Scale-Structure systematics. In Section 7.2, factor 1 could result in false fluctuations of galaxies that need to be corrected with weights. Since target selection defines color cut to select LRGs, this shift is most evident around the color selection boundaries. This effect needs to be carefully corrected since its impact on different magnitudes of LRGs may vary significantly. Factor 2 is purely due to false binning and should not be corrected at all. Ideally, systematics-weighted galaxies should still have a positive trend against the SFD $E(B-V)$ map, as we should not correct the trend that stems from factor 2.
@@
Our study on galaxy brightness in Section 7.1 and psfsize in Section 7.3 challenges Assumption 2. We show that the imaging systematics trend for DESI LRGs changes significantly when we apply a cut on $z$-band magnitude. The mechanism behind this is well-understood: Faint galaxies have more photometric scattering than bright galaxies so they are affected by the variation of galaxy depth and seeing (psfsize) to a larger extent. If two galaxies of different magnitudes have the same configuration in a survey property map, they could be affected by imaging systematics differently. However, the current survey-property-map-based approach will assign them the same “imaging systematics weight”. Applying the same correction for galaxies with different brightness will result in an over-correction in brighter galaxies and an under-correction in fainter galaxies. In both cases, they introduce an artificial signal when computing the sample’s power spectrum.
@@
One way to mitigate this issue is to separate galaxy samples into subsamples with different magnitude ranges. Then, we compute imaging systematics weights independently for each subsample. However, with a lower density for each subsample, the computation of imaging systematics weight would be prone to fitting the noise of each subsample. Thus, such a step needs to be carefully validated.
@@
Another way to deal with the issue is through forward modeling tools like Obiwan. Forward modeling can produce randoms with detailed scattering features for each source at any location. We can use such randoms to correct for magnitude-dependent imaging systematics. The randoms produced from forward modeling can predict density fluctuations at any color and brightness. Compared with making such predictions with real data, there are many advantages of predicting this density fluctuation with forward modeling randoms like Obiwan LRGs. These randoms do not contain clustering signals, ensuring recovery of imaging systematics trends solely from uniform random fluctuations. They also lack cosmic variance: Their covariance with survey property maps is smaller than real galaxies under the assumption of the same number density. Moreover, they can reach a number density larger than real galaxies with more simulations. These advantages would help make the predictions precise and free from over-fitting the true clustering signal.
Due to the scope of this work, we did not perform clustering measurements. Current Obiwan LRGs does not have a large enough density to be used as magnitude-dependent randoms. In Appendix A.4, we explore various methods to enhance the computational efficiency of generating Obiwan randoms.
7.6 Implications for cosmological analysis
For the robustness of cosmological analyses, accurately estimating and mitigating the effects of imaging systematics is crucial and will become increasingly challenging as statistical uncertainties decrease. Our work has discussed several phenomena that may go under the radar of many systematics mitigation methods: A sample could have a null systematics trend over a surveyed property in terms of number density, but its redshift distribution, galaxy bias, etc., could still change. [90] shows that galaxies’ morphology correlates with the Large-Scale-Structure environment they stay in, indicating that the morphology dependence possibly links to bias differences. The changes in the galaxy properties and thus (likely) galaxy bias parameters may complicate measurements of structure growth that combine galaxy auto correlations and galaxy-lensing cross-correlations (e.g., [91]), as these typically assume the same effective galaxy bias parameters in the modeling to recover the amplitude of matter clustering. The variation in galaxy bias is particularly problematic for primordial non-Gaussianity measurements. Many survey properties vary considerably on large scales (e.g., $E(B-V)$). When the galaxy bias shifts with these survey properties, it effectively introduces a scale-dependent bias that could degenerate with the primordial non-Gaussianity signal. [92] found that the need to include many maps in the modeling of imaging systematics inflated the statistical uncertainty by more than a factor of two. By correctly modeling how the intrinsic properties of galaxies change with survey property maps, we can reduce this statistical uncertainty induced by imaging systematics.
@@ 8 Conclusions
  • We find shape bias in the COSMOS Deep catalog. We develop a procedure to eliminate this shape bias successfully, see Section 5.1 and Appendix C.
@@
  • We see imprints of systematics in extinction maps based on dust emission (e.g. CIB contamination, variation in dust properties that changes the optical-infrared relation). We find significant trends between the DESI LRGs density on the sky and the SFD $E(B-V)$ map [66], which are not predicted by Obiwan. These trends change significantly when we test against alternative $E(B-V)$ maps and are nearly removed when compared to the stellar-reddening-based $E(B-V)$ map [70]. Since CIB is correlated with galaxy clustering, we caution against any analyses of DESI LRGs eliminating this trend, as this would result in removing real large-scale structures. Meanwhile, some extent of the trends with $E(B-V)$ is real and related to the impact on the imaging depth (photometric scattering), and is predicted by our forward modeling.
  • We find that Obiwan predicts the trends with psfdepth that we observe in the real data, which are similarly strong in each of the optical imaging bands. While the overall trends are predicted, statistically, they are not consistent in the SGC. The slight mismatch we see is due to cosmic variance in the truth input, which has a larger impact on the deepest data in the SGC. The mismatch in the color and magnitude distribution of Obiwan LRGs and DESI LRGs is the main source of the systematics trend deviation in this region. We further demonstrate that this issue can be greatly mitigated by applying a ‘color matching’ weight to Obiwan LRGs to match its color distribution with that of DESI LRGs. The details are discussed in Section 7.1.
  • We find that the trends with imaging depth depend strongly on the $z$-band magnitude. When selecting the data to have $z<20.5$, only  1% level variations are found in the observed data with the imaging depth, and Obiwan predicts no significant trend. However, with selection cut $z>20.5$, up to 20% trends are observed in the real data and predicted by Obiwan. This suggests it is important to properly model the color/brightness dependence of systematic trends in the DESI LRGs target sample when producing any corrections for use in galaxy clustering measurements. Forward modeling randoms like DESI LRGs can be a powerful catalog to correct for imaging systematics that varies with intrinsic galaxy brightness.
  • We find signs of non-simulated density variation against psfsize, which is especially prominent for extended galaxies. This is likely due to a bias in background subtraction that is correlated with psfsize for each CCD. This effect results in LRGs around color selection boundaries having different properties, and the impact is also seen in LRGs after spectroscopic redshift cut.
@@
  • By comparing systematic trends in real galaxies with those in synthetic galaxies generated by forward modeling tools like Obiwan, we can detect potential unknown imaging systematics and explore relevant resources to identify their root causes. Replicating imaging systematics in real galaxies through forward modeling tools like Obiwan is challenging work. A perfect match between real and synthetic galaxies is only possible when all potential contaminations are properly considered in the image simulation settings. However, a mismatch of density fluctuations in the synthetic galaxies also provides us valuable insights into what unknown systematics could possibly be missed. Identifying missing factors is vital for proper imaging systematics mitigation.
@@
We produce the Obiwan LRGs catalog that traces imprints of DESI LRGs in the real images and can be used for future studies. The product is currently available internally for DESI members  [^8] https://data.desi.lbl.gov/desi/survey/catalogs/image_simulations/LRG/. The data will be uploaded to zenodo before publication [^9] https://zenodo.org/uploads/11260267. In the future, the catalogs used in this analysis will be made public along with the Data Release 1 101010 https://data.desi.lbl.gov/doc/releases/.
@@ Acknowledgements
We thank Alex Krolewski and the anonymous reviewer for reading and providing feedback on this work.